Mass density of concrete is 2.3x103kg/m3. Calculate the height of the tallest cylindrical concrete pillar that will not collapse under its own weight. The compression strength of concrete is 2.0x107N/m2


1
Expert's answer
2021-07-02T09:38:58-0400

"mg=ES,"

"\rho Vg=ES,"

"\rho Shg=ES,"

"\rho hg=E,"

"h=\frac E{\rho g},"

"h=\frac{2\cdot 10^7}{2.3\cdot 10^3\cdot9.8}=887~m."